# Optimizing ML Concurrent Computation and Communication with GPU DMA Engines

### v1

@@ Abstract
Concurrent computation and communication (C3) is a pervasive paradigm in ML and other domains, making its performance optimization crucial. In this paper, we carefully characterize C3 in ML on GPUs, which are most widely deployed for ML training and inference. We observe that while C3 leads to performance uplifts, the uplifts are far lower than ideal speedups (serial computation and communication versus maximum of computation or communication; all times from isolated executions). C3 on average achieves only 21% of ideal speedup, this is due to known challenges of compute and memory interference between concurrent GPU kernels (that is, sharing of GPU’s compute units, caches and HBM).
@@
To achieve this, we first create a taxonomy of C3 and provide detailed characterization of C3 anchored on this taxonomy. Given our ML focus, we study C3 manifestations wherein the computation kernel is a matrix-matrix multiplication (or GEMM) kernel and the communication kernel is a collective operation such as an all-gather among multiple GPUs. By analyzing C3 scenarios from the training of LLaMA-70B and LLaMA-405B [4] models, along with some synthetic C3 scenarios, we provide broad coverage for our proposed taxonomy.
@@
Next, using detailed experiments, we determine the compute and memory bandwidth requirements of isolated executions of the computation and communication kernels under study. This enables us to get an assessment of ideal performance uplift via concurrency. We observe that, when it comes to compute needs, on a state-of-art GPU such as AMD Instinct™ MI300X which has ample compute units (or GPU compute cores), communication kernels need only up to 10-20% of available units making the rest available to concurrent computation kernels (Figure 1, left). GEMMs, as expected, manifest a spectrum of behaviors, some resilient to losing compute units to concurrent communication while the rest being sensitive. On the memory front, except for memory bandwidth-bound GEMM kernels, we observe that the bandwidth needs of compute-bound GEMMs and communication kernels can be met in tandem by the high memory bandwidth made available by MI300X GPUs. Overall, for C3 scenarios under study, compared to a baseline which serializes computation and communication, if the shorter of the computation or communication kernels is completely hidden, ideal speedups of 1.6 (average) and 2 (maximum) are possible.
@@
However, here we observe that of this ideal speedup, on average only 21% (1.13) is realized. This is not unexpected, as concurrent GPU kernels cause mutual interference with each other as they share compute cores, caches, high bandwidth memory (HBM) bandwidth and more. That is, available resources to a given GPU kernel are lower in presence of other kernels than in isolation. In fact, prior work [5] has observed that due to such interference, C3 can even lead to slowdowns and loss of ML throughput.
@@
To bridge the above performance gap, we first evaluate two strategies: schedule prioritization and careful partitioning of GPU compute units. Specifically, using our isolated execution analysis above, we observe that prioritizing scheduling of kernels with lower resource requirements prevents starvation and leads to overall better performance. Similarly, we also demonstrate that the resource partitioning available on MI300X GPUs wherein certain compute units on GPUs can be exclusively marked for a given stream of work can be harnessed to push C3 performance. Overall, using these strategies we can push performance attained with C3 to an average of about 42% of available ideal speedup. Additionally, we also provide heuristics that can guide a runtime in employing these strategies.
@@
Finally, to further enhance C3 performance, we tackle the interference incurred with C3 manifestations today. That is, for compute interference, instead of splitting available compute units among compute and communication kernels (Figure 1, left), we harness existing direct memory access (DMA) engines on MI300X GPU and offload communication to them making all compute units available to concurrent GEMM computations (Figure 1, right). Given the placement of DMA engines in MI300X GPU (Section II-A), this has the added advantage of lowering interference in subset of caches as well. To this end, we build ConCCL proof-of-concepts (PoCs) for ML collectives such as all-gather and all-to-all111Note that, as GPU DMA engines do not support arithmetic operations, we do not consider offloading all-reduce kernels to DMAs.. We demonstrate that our simple PoCs are at par with existing communication libraries for bandwidth-bound scenarios while also reducing interference in C3, leading to attainment of on average 72% of ideal speedup, up to a max of 1.67 speedup. Overall, our work makes a strong case for continued investment and betterment of GPU DMA engines and we conclude with a discussion of further GPU enhancements to support C3 efficiently.
@@
To attain further performance uplifts, unlike current state-of-art communication libraries which offload communication to GPU compute-units, we build ConCCL (Concurrent Communication CoLlectives) proof-of-concepts, which offload communication to DMA engines on GPUs.
@@ II-A AMD Instinct MI300X GPU and Compute Orchestration
Large-scale ML (the focus of this work) often employs multiple GPUs in tandem. In this work, we focus on the AMD MI300X Infinity Platform comprised of a 8x MI300X node with a fully-connected topology. Each MI300X connects to seven other MI300X GPUs using AMD Infinity Fabric™ [8] bi-directional links (each uni-directional bandwidth of 64GB/s/link).
@@ II-B Direct Memory Access (DMA) Engines in GPUs
As discussed above, each MI300X is comprised of 14 SDMA copy engines available for transfers between address-space shared GPUs. We depict the steps involved for users to use these DMA engines in Figure 3. Using either heterogeneous-computing interface for portability (HIP) [10] or heterogeneous system architecture (HSA) [9] runtime API calls, at user-level, a programmer requests a single data transfer to be done using SDMA engines. Under the hood, this causes the runtime on CPU to place a command packet ( ) in the DMA queue placed in system memory for a specific GPU (either source or destination GPU for the transfer). The DMA engine of the GPU polls this queue and pulls the command from the queue ( ) and processes it ( ). Once decoded, the DMA engine issues necessary reads/writes from HBM memory of source/destination GPU respectively to complete the transfer ( ).
@@ II-C ML Operators and C3 in ML
As large-scale ML training and inference increasingly relies on GPU clusters, GPUs need to support both efficient computation and efficient communication. Furthermore, there are several ML algorithmic choices that lead to concurrent computation and communication. C3 entails a pair of computation kernel and communication kernel that have no data dependencies and as such as be scheduled concurrently on a given GPU.
Examples of such algorithmic choices include data-parallelism [1], FSDP [2], nano-batching [3] and more. In data-parallelism, back propagation phase of ML training comprises matrix-matrix multiplication gradient kernels (both input and weight gradients) which can be scheduled concurrently with reduction (communication) of weight gradients across GPUs for intra-layer concurrency (input gradient) or inter-layer concurrency (weight gradient of previous layer). Similarly techniques like FSDP, gather model weights for a given layer on a GPU (communication), while performing computations of previous layers. Finally, optimizations like nano-batching [3] break down a single batch of inputs into nano-batches opening up opportunity for compute and communication kernels from different nano-batches to be co-scheduled on a GPU. Overall, given the ample prevalence of C3 in ML, characterization of and optimization for C3 on GPUs is important and hence is the focus of this work.
@@ III C3 Taxonomy
As discussed in Section II-C, the focus of our work is C3 manifestations in ML. More specifically, while there exist a variety of operators/computations in ML, in this work we focus on two primary operators, namely, general matrix-matrix multiplications (or GEMM) kernels and communication kernels for they contribute to majority of the ML execution time in both training and inference scenarios across a spectrum of models [11]. Based on this, we depict our proposed C3 taxonomy in Figure 4. Specifically, we consider three key types of C3: G-long, C-long and GC-equal.
As their names suggest, G-long is a C3 manifestation where the computation kernel (or GEMM in this case) is more dominant (>%115) when comparing their isolated execution times. Similarly, C-long is a C3 manifestation where communication kernel is more dominant (>%115) in terms of execution time. Finally, GC-equal is a C3 manifestation where both kernels are comparable (within 15% of each other).
Even within this framework, we further observe that, the relative magnitude ( in Figure 4) of computation and communication kernels with regards to each other has an effect on expected interference incurred and as such performance attained. That is, when two kernels run concurrently on a GPU, they interfere with each other in both compute (splitting of compute units) and memory sub-system (caches, HBM). As such, we consider this part of our taxonomy and we vary the relative magnitudes of C3 scenarios under study in our work.
Finally, not all GEMM or communication kernels are the same. For GEMMs, we consider two broad categories of compute-bound and memory-bound ( in Figure 4) kernels. We define a kernel to be compute-bound if its measured op-to-byte ratio is larger than machine op-to-byte as calculated from peak compute and memory throughput of underlying processor (kernel is memory-bound otherwise). For communication, we consider two types of commonly used ML multi-GPU collective kernels - all-gather and all-to-all. Similar to GEMM kernels, such ML collective kernels, depending on communication size, can be latency-bound or bandwidth-bound ( ). We classify a communication kernel with its associated size as latency-bound if the kernel latency at/before this size does not increase commensurate to size. We further discuss the specific sizes and C3 manifestations we consider to cover this taxonomy, the source (ML model) for these manifestations and rationale in Section IV-A.
@@ IV-A2 C3 Manifestations Under Study
We discuss the GEMM kernels we study in this work, their sizes and their sources in Table I. As listed, we source our GEMM sizes from training of LLaMA-70B and LLaMA-405B [4] models assuming a total of 8192 tokens in a batch. Further, for ease of reading, we tag these GEMM sizes/kernels as compute-bound (cb) or memory-bound (mb). Recall that, we define a kernel to be compute-bound if its measured op-to-byte ratio is larger than machine op-to-byte as calculated from peak compute and memory throughput of underlying processor (kernel is memory-bound otherwise).
@@
We also list the taxonomy for each C3 in Table II. Notice that, we have more of G-long scenarios than C-long scenarios than GC-equal scenarios. This is so, as shown in Table II, majority of C3 manifestations in LLaMA models today are of G-long type and we consciously wanted to limit the synthetic scenarios we added. That said, we provide all of our results grouped by C3 type to understand the performance for each of these different types.
@@ IV-B3 Ideal speedup projection
Based on above isolated execution analysis of compute and memory needs, we project in Figure 7 the ideal speedup possible for C3 scenarios under study. For this analysis, we assume that the best speedup possible is when the smaller of the two kernels (computation or communication) happens completely in the shadow of the other. As depicted, wide variety of speedups are possible (1.1 to close to 2) and this is dictated by the relative magnitudes of these kernels (Section III). However, based on compute/memory needs analysis, due to compute/memory interference not all of this ideal performance can be attained.
@@ IV-C C3 Characterization
We present characterization of C3 for different scenarios under study in Figure 8. In this figure we refer to baseline C3 performance as c3_base. Note that, with concurrent scheduling, we schedule GEMM kernel first (as scheduling both the kernels at the exact same time is unrealistic, we minimized the scheduling delay between the two through code optimizations) in c3_base executions.
In Figure 8, we present average speedups for different groupings of C3 scenarios (collectives, and taxonomy). We also show in the graph the ideal speedups possible. The figure shows that c3_base speedups range from no speedups to up to 1.3 speedups. On average, c3_base attains 1.13 speedup which is about 21% of ideal speedup. Note that, this is not unexpected. In fact, prior work [5] observed slowdowns with C3 due to mutual interference of concurrent kernels.
@@ V-A Schedule Prioritization
Note that, thus far, for concurrent scheduling, we chose to schedule GEMM kernel first. But based off our compute needs analysis, we observe that communication kernels have far lower CU needs as compared to GEMM kernels (Section IV-B1). As such, if GEMM kernels are scheduled first, the internal GPU scheduler can in some cases allocate majority of CUs to GEMM kernel leading to potential starvation. In contrast, if we employ, schedule prioritization, that is schedule communication kernel first, later scheduled GEMM kernel will definitely have compute resources to be allocated.
We perform this schedule prioritization, referred to Figure 8 as c3_sp and show the speedups attained. Across all C3 types and collectives, we observe that schedule prioritization improves the speedups attained for C3. Overall, instead of 0-13% of ideal speedup, for all-to-all, c3_sp attains 27-46% instead. Similarly, instead of 24-46% of ideal speedup, for all-gather, c3_sp attains 38-67% of ideal speedup. Overall, on average, c3_sp helps attain 42% of ideal speedup (up from 21% of c3_base).
@@ V-B Resource Partitioning
We show this in Figure 8 as c3_rp which adds resource partitioning (rp) to c3_base. We sweep all possible powers-of-two CU allocations for communication kernels (and consequently take CUs away from GEMM kernel) and plot the best performing one as c3_rp. As depicted, c3_rp delivers performance improvements over c3_base which is almost the performance delivered by c3_sp (41% of ideal speedup). Note, that, adding resource partitioning to c3_sp, shown as c3_sp_rp, did not improve performance any further.
@@ V-C Runtime Heuristics for Improved C3 Performance
Heuristic for Resource Partitioning: For resource partitioning across concurrent kernels, we describe a runtime heuristic to allocate CUs across concurrent kernels. To do so, we picked one memory-bound GEMM kernel, one compute-bound GEMM kernel, and two all-gather and all-to-all collective sizes (latency-bound, bandwidth-bound) and used the compute need analysis done in Section IV-B1 to build a lookup table of potential slowdowns when the kernel loses given number of CUs. Note that, for a given GPU this is to be done once.

### v2

@@ Abstract
Concurrent computation and communication (C3) is a pervasive paradigm in ML and other domains, making its performance optimization crucial. In this paper, we carefully characterize C3 in ML on GPUs, which are most widely deployed for ML training and inference. We observe that while C3 leads to performance uplifts, the uplifts are far lower than ideal speedups (serial computation and communication versus maximum of computation or communication; all times from isolated executions). That is, C3 on average achieves only 21% of ideal speedup. This is so, due to known challenges of compute and memory interference between concurrent GPU kernels (that is, sharing of GPU’s compute units, caches and HBM).
@@
To achieve this, we first create a taxonomy of C3 and provide a detailed characterization of C3 anchored on this taxonomy. Given our ML focus, we study C3 manifestations wherein the computation kernel is a matrix-matrix multiplication (or GEMM) kernel and the communication kernel is a collective operation such as an all-gather among multiple GPUs. By analyzing C3 scenarios from the training of the LLaMA-70B and LLaMA-405B [4] models, along with some synthetic C3 scenarios, we provide broad coverage for our proposed taxonomy.
@@
Next, using detailed experiments, we determine the compute and memory bandwidth requirements of isolated executions of the computation and communication kernels under study. This enables us to get an assessment of ideal performance uplift via concurrency. We observe that, when it comes to compute needs, on a state-of-the-art GPU such as AMD Instinct™ MI300X which has ample compute units (or GPU compute cores), communication kernels need only up to 10-20% of available units, making the rest available to concurrent computation kernels (Figure 1, left). GEMMs, as expected, manifest a spectrum of behaviors, some resilient to losing compute units to concurrent communication while the rest are sensitive. On the memory front, except for memory bandwidth-bound GEMM kernels, we observe that the bandwidth needs of compute-bound GEMMs and communication kernels can be met in tandem by the high memory bandwidth made available by MI300X GPUs. Overall, for C3 scenarios under study, compared to a baseline which serializes computation and communication, if the shorter of the computation or communication kernels is completely hidden, ideal speedups of 1.6 (average) and 2 (maximum) are possible.
@@
However, here we observe that of this ideal speedup, on average only 21% (1.13) is realized. This is not unexpected, as concurrent GPU kernels cause mutual interference with each other as they share compute cores, caches, high bandwidth memory (HBM) bandwidth and more. That is, the available resources for a given GPU kernel are lower in the presence of other kernels than in isolation. In fact, prior work [5] has observed that due to such interference, C3 can even lead to slowdowns and loss of ML throughput.
@@
To bridge the above performance gap, we first evaluate two strategies: schedule prioritization and careful partitioning of GPU compute units. Specifically, using our isolated execution analysis above, we observe that prioritizing scheduling of kernels with lower resource requirements prevents starvation and leads to overall better performance. Similarly, we also demonstrate that the resource partitioning available on MI300X GPUs, wherein certain compute units on GPUs can be exclusively marked for a given stream of work, can be harnessed to push C3 performance. Overall, using these strategies, we can push the performance achieved with C3 to an average of about 42% of available ideal speedup. Additionally, we also provide heuristics that can guide a runtime in employing these strategies.
@@
Finally, to further enhance C3 performance, we tackle the interference incurred with C3 manifestations today. That is, for compute interference, instead of splitting available compute units among compute and communication kernels (Figure 1, left), we harness existing direct memory access (DMA) engines on the MI300X GPU and offload communication to them making all compute units available for concurrent GEMM computations (Figure 1, right). Given the placement of DMA engines in MI300X GPU (Section II-A), this has the added advantage of lowering interference in subset of caches as well. To this end, we build concurrent communication collectives (ConCCL) proof-of-concepts (PoCs) for ML collectives such as all-gather and all-to-all111Note that, as GPU DMA engines do not support arithmetic operations, we do not consider offloading all-reduce kernels to DMAs.. We demonstrate that our simple POCs are on par with existing communication libraries for bandwidth-bound scenarios while also reducing interference in C3, leading to attainment of on average 72% of ideal speedup, up to a maximum of 1.67 speedup. Overall, our work makes a strong case for continued investment and betterment of GPU DMA engines, and we conclude with a discussion of further GPU enhancements to support C3 efficiently.
@@
To attain further performance uplifts, unlike current state-of-art communication libraries which offload communication to GPU compute-units, we build concurrent communication collectives (ConCCL) proof-of-concepts, which offload communication to DMA engines on GPUs.
@@ II-A AMD Instinct MI300X GPU and Compute Orchestration
Large-scale ML (the focus of this work) often employs multiple GPUs in tandem. In this work, we focus on the AMD MI300X Infinity Platform comprised of a 8x MI300X node with a fully-connected topology. Each MI300X connects to seven other MI300X GPUs using AMD Infinity Fabric™ [8] bi-directional links (each with uni-directional bandwidth of 64GB/s/link).
@@ II-B Direct Memory Access (DMA) Engines in GPUs
As discussed above, each MI300X is comprised of 14 SDMA copy engines available for transfers between address-space shared GPUs. We depict the steps involved for users to use these DMA engines in Figure 3. Using either heterogeneous-computing interface for portability (HIP) [10] or heterogeneous system architecture (HSA) [9] runtime API calls, at user-level, a programmer requests a single data transfer to be done using SDMA engines. Under the hood, this causes the runtime on CPU to place a command packet ( ) in the DMA queue placed in system memory for a specific GPU (either source or destination GPU for the transfer). The GPU DMA engine gets notified and fetches the command from the queue ( ) and processes it ( ). Once decoded, the DMA engine issues necessary reads/writes from/to HBM memory of source/destination GPU respectively to complete the transfer ( ).
@@ II-C ML Operators and C3 in ML
As large-scale ML training and inference increasingly rely on GPU clusters, GPUs need to support both efficient computation and efficient communication. Furthermore, there are several algorithmic choices in ML that lead to concurrent computation and communication. C3 entails a pair of computation kernel and communication kernel that have no data dependencies and as such as be scheduled concurrently on a given GPU.
Examples of such algorithmic choices include data-parallelism [1], FSDP [2], nano-batching [3] and more. In data parallelism, the backpropagation phase of ML training comprises matrix-matrix multiplication gradient kernels (both input and weight gradients), which can be scheduled concurrently with reduction (communication) of weight gradients across GPUs for intra-layer concurrency (input gradient) or inter-layer concurrency (weight gradient of previous layer). Similarly, techniques like FSDP, gather model weights for a given layer on a GPU (communication), while performing computations of previous layers. Finally, optimizations like nano-batching [3] break down a single batch of inputs into nano-batches opening up opportunity for compute and communication kernels from different nano-batches to be co-scheduled on a GPU. Overall, given the ample prevalence of C3 in ML, characterization and optimization of C3 on GPUs is important and hence the focus of this work.
@@ III C3 Taxonomy
As discussed in Section II-C, the focus of our work is on the manifestations of C3 in ML. More specifically, while there exist a variety of operators/computations in ML, in this work we focus on two primary operators, namely general matrix-matrix multiplications (or GEMM) kernels and communication kernels, as they contribute to the majority of the ML execution time in both training and inference scenarios across a spectrum of models [11]. Based on this, we depict our proposed C3 taxonomy in Figure 4. Specifically, we consider three key types of C3: G-long, C-long and GC-equal.
We use execution times in isolation for our taxonomy. Thus, G-long is a C3 manifestation where GEMM time in isolation is >115% of communication time. Similarly, C-long implies communication time in isolation is >115% of GEMM time. Finally, GC-equal is a C3 manifestation where both kernels are comparable (within 15% of each other).
Even within this framework, we further observe that the relative magnitude ( in Figure 4) of computation and communication kernels with respect to each other has an effect on expected interference incurred and as such performance attained. That is, when two kernels run concurrently on a GPU, they interfere with each other in both compute (splitting of compute units) and memory subsystem (caches, HBM). As such, we consider this part of our taxonomy and vary the relative magnitudes of the C3 scenarios under study in our work.
Finally, not all GEMM or communication kernels are the same. For GEMMs, we consider two broad categories of compute-bound and memory-bound ( in Figure 4) kernels. We define a kernel to be compute-bound if its measured op-to-byte ratio is larger than machine op-to-byte as calculated from peak compute and memory throughput of underlying processor (kernel is memory-bound otherwise). For communication, we consider two types of commonly used multi-GPU ML collective kernels - all-gather and all-to-all. Similar to GEMM kernels, such ML collective kernels, depending on communication size, can be latency-bound or bandwidth-bound ( ). We classify a communication kernel with its associated size as latency-bound if the kernel latency at/before this size does not increase commensurate to size. We further discuss the specific sizes and C3 manifestations that we consider to cover this taxonomy, the source (ML model) for these manifestations and rationale in Section IV-A.
@@ IV-A2 C3 Manifestations Under Study
We discuss the GEMM kernels we study in this work, their sizes and their sources in Table I. As listed, we source our GEMM sizes from training of LLaMA-70B and LLaMA-405B [4] models processing 8192 tokens (i.e., product of input length and batch-size) in a given iteration. Further, for ease of reading, we tag these GEMM sizes/kernels as compute-bound (cb) or memory-bound (mb). Recall that, we define a kernel to be compute-bound if its measured op-to-byte ratio is larger than machine op-to-byte as calculated from peak compute and memory throughput of underlying processor (kernel is memory-bound otherwise).
@@
We also list the taxonomy for each C3 in Table II. Notice that we have more G-long scenarios than C-long scenarios than GC-equal scenarios. This is so, as shown in Table II, majority of C3 manifestations in LLaMA models today are of G-long type and we consciously wanted to limit the synthetic scenarios we added. However, we provide all our results grouped by C3 type to understand the performance for each of these different types.
@@ IV-B3 Ideal speedup projection
Based on above isolated execution analysis of compute and memory needs, we project in Figure 7 the ideal speedup possible for C3 scenarios under study. For this analysis, we assume that the best speedup possible is when the smaller of the two kernels (computation or communication) happens completely in the shadow of the other. As depicted, wide variety of speedups are possible (1.1 to close to 2) and this is largely dictated by relative magnitudes of these kernels (Section III). However, based on compute/memory needs analysis, due to compute/memory interference not all of this ideal performance can be attained.
Further, note that, based on runtime decisions or GPU-GPU execution variation or kernel-size/type, different degrees of overlap can manifest, resulting in different ideal speedups. To precisely tackle this we design C3 taxonomy so that we can consider such varied manifestations and analyze our proposals across these different C3 manifestations.
@@ IV-C C3 Characterization
We present characterization of C3 for different scenarios under study in Figure 8. In this figure we refer to baseline C3 performance as c3_base. Note that, with concurrent scheduling, we schedule GEMM kernel first (as scheduling both the kernels at the exact same time is challenging, we minimized the scheduling delay between the two through code optimizations) in c3_base executions.
In Figure 8, we present average speedups for different groupings of C3 scenarios (collectives, and taxonomy). We also show in the graph the ideal speedups possible (marked at the top of the graph). The figure shows that c3_base speedups range from no speedups to up to 1.3 speedups. On average, c3_base attains 1.13 speedup which is about 21% of ideal speedup. Note that, this is not unexpected. In fact, prior work [5] observed slowdowns with C3 due to mutual interference of concurrent kernels.
@@ V-A Schedule Prioritization
Note that, thus far, for concurrent scheduling, we chose to schedule GEMM kernel first. But based off our compute needs analysis, we observe that communication kernels have far lower CU needs as compared to GEMM kernels (Section IV-B1). As such, if GEMM kernels are scheduled first, the internal GPU scheduler can in some cases allocate majority of CUs to GEMM kernel leading to potential starvation. In contrast, if we employ, schedule prioritization, that is, schedule communication kernel first, later scheduled GEMM kernel will definitely have compute resources to be allocated. Overall, the key insight of schedule prioritization is that prioritizing or providing quality of service for kernel with smaller and complimentary resource requirement helps concurrent performance.
To realize schedule prioritization, from CPU-side, we first schedule communication kernel in a stream and then immediately after, schedule GEMM kernel in concurrent stream. We present the results of schedule prioritization, referred in Figure 8 as c3_sp and show the speedups attained. Across all C3 types and collectives, we observe that schedule prioritization improves the speedups attained for C3. Overall, instead of 0-13% of ideal speedup, for all-to-all, c3_sp attains 27-46% instead. Similarly, instead of 24-46% of ideal speedup, for all-gather, c3_sp attains 38-67% of ideal speedup. Overall, on average, c3_sp helps attain 42% of ideal speedup (up from 21% of c3_base).
@@ V-B Resource Partitioning
We show this in Figure 8 as c3_rp which adds resource partitioning (rp) to c3_base. We sweep all possible powers-of-two CU allocations for communication kernels (and consequently take CUs away from GEMM kernel) and plot the best performing one as c3_rp. As depicted, c3_rp delivers performance improvements over c3_base which is almost the performance delivered by c3_sp (41% of ideal speedup). Note, that, adding resource partitioning to c3_sp, depicted as c3_sp_rp, did not improve performance any further.
@@ V-C Runtime Heuristics for Improved C3 Performance
Heuristic for Resource Partitioning: For resource partitioning across concurrent kernels, we describe a runtime heuristic to allocate CUs across concurrent kernels. To do so, we picked one memory-bound GEMM kernel, one compute-bound GEMM kernel, and two all-gather and all-to-all collective sizes (latency-bound, bandwidth-bound) and used the compute need analysis done in Section IV-B1 to build a lookup table of potential slowdowns when the kernel loses a given number of CUs. Note that, for a given GPU this is to be done once.